Hi! Please help me solve this age related problem: When the father was 36 years old, his son was 8 years old. If the father is now 3 times as old as his son, how old is the son? The only thing I do is the listing method, but do you have any faster strategies for this? Thanks a lot!
Answer by ankor@dixie-net.com(22740)


When the father was 36 years old, his son was 8 years old.
From this statement, we know their age difference
f = s + 28
Write an equation for the statement
"If the father is now 3 times as old as his son, how old is the son?"
f = 3s
replace f with (s+28)
s + 28 = 3s
28 = 3s - s
28 = 2s
s = 28/2
s = 14 yrs is the son